Mrs. Wilson is five times as old as her daughter Kori. In five years Mrs. Wilson will be 3 times as old as Kori. How old is each at the present time?
I know that Mrs. Wilson is 25 and Kori is 5 (by process of elimination) but don't know how to set up the equation and how to solve.
Thanks for any help!
This question is not from a textbook......the teacher handed it out on a typed sheet.

Answer by stanbon(75887)

Mrs. Wilson is five times as old as her daughter Kori. In five years Mrs. Wilson will be 3 times as old as Kori. How old is each at the present time?
-------------------
NOW DATA:
Let daughter's age be x yrs ; Mother's age is 5x yrs.
------------------------------
5 yrs. from NOW DATA:
Daughter will be x+5 yrs ; Mother's age will be 5x+5 yrs.
-----------------------------
EQUATION:
In 5 yrs. Mother will be 3 times as old as daughter:
5x+5 = 3(x+5)
5x+5 = 3x+15
2x = 10
x = 5 (daughter's age NOW)
5x = 15 (Mother's age NOW)
============================
Cheers,
Stan H.